At the beginning of the period, the Cutting Department budgeted direct labor of $128,000, direct materials of $166,000 and fixed factory overhead of $10,300 for 8,000 hours of production. The department actually completed 11,900 hours of production. What is the appropriate total budget for the department, assuming it uses flexible budgeting?

Round your final answer to the nearest dollar. Do not round interim calculations.

a.$447,625

b.$304,300

c.$309,321

d.$452,646
Answer #1

Option A is correct

direct material =166000

direct labour = 128000

fixed over heads = 10300

value of excess hours of production = 143325

total budget for the department = 447625

Working Notes :

calculation of value for  Actual hours

Rate per hour = (direct materials + direct labour) / budgeted hours

= (128000 + 166000) / 8000

= 294000 / 8000

= 36.75

value of excess hours of production using rate per hour = (11900 - 8000) * 36.75

= 143325

Thank you so much

comment for any help